# Copyright Statement This paper has been accepted for publication in the IEEE ICBASE 2024 conference

### v1

@@ Abstract
Lung cancer remains one of the most prevalent and lethal forms of cancer worldwide, accounting for approximately 12% of all new cancer diagnoses and nearly 20% of all cancer deaths annually. It is primarily classified into two major types: non-small cell lung cancer (NSCLC), which represents about 82% of cases, and small cell lung cancer (SCLC), which is less common but more aggressive. Despite advancements in early detection and treatment, the prognosis for lung cancer patients remains poor, with a five-year survival rate of approximately 25%, underscoring the critical need for continued research and innovation in this field [1].
@@
Stage III non-small cell lung cancer (NSCLC) is a complex disease, where accurate prognostic evaluation is key to personalized treatment. Detecting prognostic biomarkers can guide therapeutic decisions, potentially benefiting patients eligible for immunotherapy or targeted therapies. Recent studies, including a review on stage III NSCLC management, highlight the potential of biomarker-driven approaches to refine patient selection, improve outcomes, and advance personalized treatment strategies in this challenging lung cancer subset [3]. Given the poor prognosis of stage III NSCLC, reliable biomarkers could significantly enhance treatment effectiveness and survival rates [4].
@@
The rapid expansion of the bioinformatics field such as gene expression profiling has also led to a growing dependence on machine learning techniques[7] for diagnosing and predicting complicated diseases with their biomarkers. However, the biomedical data is often high-dimensional, with a large quantity of variables and limited observational data, as well as multicollinearity, which is challenging to analyze. To address this, researchers have developed various machine learning algorithms in bioinformatics, such as tree-based methods[8] as well as neural networks, which have been proven powerful for identifying complex patterns for different domains, such as computer vision[9, 10], natural language processing, and more comprehensive tasks[11, 12, 13].
@@ II-C Data Pre-Processing and Split
In this study, we split the dataset of 60 individuals into a training set and a test set, with 60% (n=36) used for training and 40% (n=24) for testing. This approach ensures that the model has sufficient data for learning while retaining enough for evaluating its performance on unseen individuals.
@@ III Methods
XGBoost[16], a tree-based machine learning method, was used in this study to identify biomarkers that contribute to stage III lung cancer. The following is the objective function:
where represents the ensemble of decision tree models, is a convex loss function that is differentiable (the gap between the actual output and the predicted output ), is the true output for the -th data point, is the corresponding predicted output, and is the penalty term[17] for each tree within the ensemble .
XGBoost optimizes[18] the target function additively, by iteratively building an ensemble of simple decision trees, or ”weak learners”, that work together to progressively reduce the objective function. During each iteration, a new tree model is added into the ensemble, further optimizing the objective function. This process can be mathematically represented as:
where represent the prediction after trees have been added, represent the prediction after trees have been added, and denote the tree added in the -th iteration.
By showcasing the potential of and XGBoost, the research[19] of Yu et al. provides a strong foundation and direction for future studies focusing on building robust and efficient classification models for various applications. We have leveraged Chang’s research findings[20], significantly enhancing our model processing efficiency through the adoption of his data handling techniques. As a tree-based model, XGBoost also has a good explainability[21], so we anticipate that it will capture complex patterns within the GDS3837 dataset to identify biomarkers associated with stage III lung cancer.

### v2

# Identification of Prognostic Biomarkers for Stage III Non-Small Cell Lung Carcinoma in Female Nonsmokers Using Machine Learning

@@ Abstract
Lung cancer is a significant health challenge globally, being among the most common and deadly cancers. It accounts for approximately 12% of all new cancer diagnoses and nearly 20% of all cancer deaths annually. It is primarily classified into two major types: non-small cell lung cancer (NSCLC), which represents about 82% of cases, and small cell lung cancer (SCLC), which is less common but more aggressive. Despite advancements in early detection and treatment, the prognosis for lung cancer patients remains poor, with a five-year survival rate of approximately 25%, underscoring the critical need for continued research and innovation in this field [1].
@@
Stage III non-small cell lung cancer is a complex disease, where accurate prognostic evaluation is key to personalized treatment. Detecting prognostic biomarkers can guide therapeutic decisions, potentially benefiting patients eligible for immunotherapy or targeted therapies. Recent studies, including a review on stage III NSCLC management, highlight the potential of biomarker-driven approaches to refine patient selection, improve outcomes, and advance personalized treatment strategies in this challenging lung cancer subset [3]. Given the poor prognosis of stage III NSCLC, reliable biomarkers could significantly enhance treatment effectiveness and survival rates [4].
@@
The rapid expansion of the bioinformatics field such as gene expression profiling has also led to a growing dependence on machine learning techniques[7] for diagnosing and predicting complicated diseases with their biomarkers. However, the biomedical data is often high-dimensional, with numerous variables and limited data points, along with the issue of multicollinearity, which often complicates analysis. To tackle these challenges, researchers have created a range of machine learning algorithms specifically for bioinformatics, such as tree-based methods[8] as well as neural networks, which have been proven powerful for identifying complex patterns for different domains, such as computer vision[9, 10], natural language processing, and more comprehensive tasks[11, 12, 13].
@@ II-C Data Pre-Processing and Split
In this study, we did a train-test split on 60 individuals, with 60% (n=36) used for training and 40% (n=24) for testing. This approach ensures that the model has sufficient data for learning while retaining enough for evaluating its performance on unseen individuals.
@@ III Methods
XGBoost[16], a tree-based machine learning method, was utilized in this study to identify biomarkers associated with stage III lung cancer. The objective function in XGBoost is designed to minimize the following expression:
In this equation, is the collection of tree models, is a convex loss function that’s differentiable, and represents the true output and the predicted output for the -th data point respectively, and is the penalty term[17] for each tree model in the collection.
XGBoost optimizes[18] the loss function described in Equation 1 through an additive approach, where a new tree is sequentially added to the model to improve the overall prediction. This iterative process can be expressed as:
where is the prediction after iterations, is the prediction from the previous iteration, represents the newly added tree in the -th iteration, and is the learning rate.
By showcasing the potential of and XGBoost, the research[19] of Yu et al. lays a solid groundwork and guidance for future research aimed at developing robust and efficient classification models for a range of applications. We have leveraged Chang’s research findings[20], significantly enhancing our model processing efficiency through the adoption of his data handling techniques. As a tree-based model, XGBoost also has a good explainability[21], so we anticipate that it will capture complex patterns within the GDS3837 dataset to identify biomarkers associated with stage III lung cancer.